OpenStudy (jaweria):

Help me Please. e^x-1-5=5.

OpenStudy (jaweria):

Helpl please

OpenStudy (anonymous):

Solve for \(e^{x-1}\) and take the natural logarithm of both sides.

OpenStudy (jaweria):

I started this problem but its giving me a wrong answer can u please explain this to me?

OpenStudy (anonymous):

Well, tell me, what do you get when you solve for \(e^{x-1}\)?

OpenStudy (jaweria):

can anyone help me here please?

OpenStudy (stamp):

\[e^x-1-5=5.\]\[e^x-6=5\]\[e^x=11\]\[ln(e^x)=ln(11)\]\[xln(e)=ln(11)\]\[x=ln(11)\]

OpenStudy (jaweria):

Sorry but the answer is wrong answer suppose to be 3.3026

OpenStudy (stamp):

@Jaweria Then you wrote it down wrong.

OpenStudy (jaweria):

ok let me fix it

OpenStudy (jaweria):

\[e ^{x-1}-5=5\]

OpenStudy (stamp):

\[e^{x-1}-5=5\]Isolate e term.\[e^{x-1}=10\]Natural log of both sides.\[lne^{x-1}=ln(10)\]Properties of logatrithms.\[(x-1)lne=ln(10)\]lne = 1.\[x-1=ln(10)\]Solve for x.\[x=ln(10)+1\]

OpenStudy (jaweria):

Sorry but I have one more question if you dont mind ?

OpenStudy (jaweria):

\[9^{n+10}+3=81\]

OpenStudy (stamp):

\[9^{n+10}+3=81\]\[9^{n+10}=81-3\]\[ln9^{n+10}=ln78\]\[n+10=ln78/ln9\]\[n=ln78/ln9-10\]

OpenStudy (jaweria):

Thanks :)

OpenStudy (stamp):

@Jaweria Properties of Logarithms http://www.andrews.edu/~calkins/math/webtexts/numb17.htm

OpenStudy (jaweria):

Thanks for this :)

OpenStudy (jaweria):

I m having trouble with this question. \[-2 \log_{5}7x=2 \]

OpenStudy (stamp):

\[-2log_57x=2\]\[log_57x=-1\]\[log_ax=y\]\[a^y=x\]\[5^{-1}=7x\]\[\frac{1}{5}=7x\]\[x=1/35\]

OpenStudy (jaweria):

Thanks

OpenStudy (jaweria):

I m working on this problem so far I got this but my answer is not matching with my professor's.